Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

logic cracks the quants...

Expert replies
by Night reader » Thu Dec 02, 2010 3:25 pm
Nora will enter a ticket lottery every day until she wins the lottery, after which she will no longer enter. If the probability that she wins the ticket lottery is 0.1 (1/10) on each of the …rest three days, what is the probability that she wins on the third day?

(A) 0.001
(B) 0.009
(C) 0.081
(D) 0.729
(E) 0.900

this problem was rated as the challenge in Gmat Hacks, though it doesn't appear difficult
My knowledge frontiers came to evolve the GMATPill's methods - the credited study means to boost the Verbal competence. I really like their videos, especially for RC, CR and SC. You do check their study methods at https://www.gmatpill.com
Join the discussion
Source: — Problem Solving |

by jaymw » Thu Dec 02, 2010 7:21 pm
Let's use some logic to solve this.

She only gets to the third day if she didn't win on the first 2 days. We can therefore multiply the probabilities of not winning for the first two days (0.9 x 0.9 = 0.81)

On the third day, she has to win to fulfill the mentioned requirement. Therefore we have to multiply 0.81 with the probability of winning (0.81 x 0.1 = 0.081).

Thus, the answer is C.
Join the discussion

by Rahul@gurome » Thu Dec 02, 2010 7:23 pm
If Nora is winning on the third day, it means she is losing on the first two days.
Probability of winning each day is 0.1.
Probability of losing each day is 1-0.1 = 0.9.
So the probability of Nora winning on the third day is 0.9*0.9*0.1 = 0.081.
The correct answer is (C).
Rahul Lakhani
Quant Expert
Gurome, Inc.
https://www.GuroMe.com
On MBA sabbatical (at ISB) for 2011-12 - will stay active as time permits
1-800-566-4043 (USA)
+91-99201 32411 (India)
Join the discussion

by Scott@TargetTestPrep » Tue Dec 12, 2017 9:39 am
Night reader wrote:Nora will enter a ticket lottery every day until she wins the lottery, after which she will no longer enter. If the probability that she wins the ticket lottery is 0.1 (1/10) on each of the …rest three days, what is the probability that she wins on the third day?

(A) 0.001
(B) 0.009
(C) 0.081
(D) 0.729
(E) 0.900
We are given that Nora’s probability of winning the lottery on each of the first 3 days is 0.1; thus, the probability of her not winning on any day is 1 - 0.1 = 0.9. We must determine the probability of her winning on day 3, which means she does not win on day 1 or on day 2.

P(winning on day 3) = 0.9 x 0.9 x 0.1 = 0.081.

Answer: C

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion